Can you help me solve this radical equation
x-2/2x+1=x+1/x

Answer by stanbon(75887)

Can you help me solve this equation
x-2/2x+1=x+1/x
------
Cross Multiply to get:
x(x-2) = (x+1)(2x+1)
---------
x^2-2x = 2x^2 + 3x + 1
---------
x^2 +5x + 1 = 0
--------
Quadratic Formula:
x = [-5 +- sqrt(25-4)]/2
-----
x = [-5 +- sqrt(21)]/2
==============
Cheers,
Stan H.
===================